# Revealing the Origin of Neutrino Masses through Displaced Shower Searches in the CMS Muon System

### v1

@@ Abstract
We study the potential to probe the origin of neutrino masses, by searching for long-lived right-handed neutrinos (RHNs) $N$ in the $B-L$ gauge model. Despite the small active-sterile mixing $|V_{\ell N}|^{2}$, RHNs are produced abundantly via SM and exotic Higgs production, as long as the Higgs mixing is sufficiently large. We reinterpret a search for displaced showers in the CMS muon system and we find that it is sensitive to parameter space at and below the canonical seesaw floor, $|V_{\ell N}|^{2}\approx 10^{-12}$ ($\ell=e$, $\tau$) for $m_{N}\approx 40$ GeV. With existing data constraining such a well-motivated scenario of neutrino mass generation, we determine the projected sensitivity at the HL-LHC, motivating dedicated searches for long-lived RHNs with decay lengths $\approx 10$ m.
@@ I Introduction
The Standard Model of particle physics cannot explain the observed neutrino masses, which have been confirmed in neutrino oscillations Davis (1994); Fukuda et al. (1998); Eguchi et al. (2003). The most popular scenario to explain the presence of small neutrino masses is through the seesaw mechanism Minkowski (1977) where heavy right-handed neutrinos (RHNs) are introduced. Through their Yukawa couplings with the Standard Model (SM) Higgs and the active neutrinos, light neutrino masses are induced. This mechanism can also generate the observed matter-antimatter asymmetry of the Universe through leptogenesis Fukugita and Yanagida (1986); Luty (1992); Davidson et al. (2008).
If the RHN masses $m_{N}$ are of the order of the electroweak (EW) scale, the active-sterile mixing strength $V_{\ell N}$ required to generate the SM neutrino mass $m_{\nu}\lesssim 0.4$ eV is $|V_{\ell N}|^{2}\sim m_{\nu}/m_{N}\lesssim 10^{-11}\times(40~{}\text{GeV}/m_{N})$. Long-lived RHNs are produced at colliders via this weak active-sterile neutrino mixing. Given the displaced signature coupled with low backgrounds, RHNs have received increased attention in phenomenological studies Bondarenko et al. (2018); Bryman and Shrock (2019); Balaji et al. (2020); Liu et al. (2022a); Abdullahi et al. (2023); Zhang and Liu (2023); Barducci et al. (2023); Liu and Deppisch (2023). Numerous experimental searches for RHNs have been carried out Chatrchyan et al. (2012); Aaij et al. (2014); Aad et al. (2015); Khachatryan et al. (2015, 2016); Cortina Gil et al. (2018); Sirunyan et al. (2018, 2019); Aad et al. (2019a); Aaij et al. (2021); Tumasyan et al. (2022); Aad et al. (2023a); Hayrapetyan et al. (2024a), with recent summaries in Bolton et al. (2020); Abdullahi et al. (2023). The resulting limits on the active-sterile mixing are $|V_{\ell N}|^{2}\lesssim 10^{-5}$ for $m_{N}\approx 10$ GeV.
While suggestive, the seesaw mechanism does not explain the origin of the RHN Majorana masses. Arguably the most direct scenario to remedy this is the $B-L$ (baryon - lepton number) gauge model Davidson (1979); Mohapatra and Marshak (1980); it generates RHN Majorana masses through the spontaneous breaking of a $U(1)_{B-L}$ gauge symmetry. The origin of heavy neutrino masses is thus analogous to that in SM EW symmetry breaking. The inclusion of three RHNs $N_{i}$, in addition to the exotic gauge boson $Z^{\prime}$ and $B-L$-breaking Higgs $\Phi$ makes this model free of quantum anomalies. The RHNs may also generate the baryon asymmetry of the universe via leptogenesis Liu and Deppisch (2023) and the $B-L$ model can be made conformally invariant. This could not only explain why the symmetry is broken near the TeV scale but could also induce EW symmetry breaking Iso et al. (2009a, b). In such a scenario, primordial black holes may explain Dark Matter Baldes and Olea-Romacho (2024). RHNs in the $B-L$ model not only interact with the SM particles via the active-sterile neutrino mixing $V_{\ell N}$, but also through the $B-L$ gauge and Higgs portals.
Thus, RHNs can also be produced via the SM Higgs, the $B-L$ Higgs as well as the SM $Z$ and $B-L$ $Z^{\prime}$. In these channels, the production is independent of the active-sterile mixing strength. Hence, the RHN can be abundantly produced if either the Higgs mixing or the $B-L$ gauge coupling are large. It is well known that enabling these portals allows testing active-sterile mixing strengths Deppisch et al. (2014); Batell et al. (2016); Deppisch et al. (2019); Accomando et al. (2018); Das et al. (2019a); Cheung et al. (2021); Chiang et al. (2019); Fileviez Pérez and Plascencia (2020); Das et al. (2019b); Liu et al. (2022b); Han et al. (2021); Das and Okada (2017); Graesser (2007); Maiezza et al. (2015); Deppisch et al. (2018); Mason (2019); Accomando et al. (2017); Gao et al. (2020); Gago et al. (2015); Jones-Pérez et al. (2020); Liu et al. (2022c); Li et al. (2024); Deppisch et al. (2023); Bernal et al. (2023). This is because the RHN decay still proceeds via $|V_{\ell N}|$ ($\ell=e,\mu,\tau$), i.e., $N\to\nu_{\ell}h,\nu_{\ell}Z^{*},\ell W^{*}$, where the off-shell SM $h$, $Z$ and $W$ decay via SM particles. The total decay width is thus suppressed by the active-sterile mixing, leading to macroscopic proper decay lengths ($m_{N}<m_{W,Z}$tre et al. (2009),
@@
Long-lived particles (LLPs), such as the RHNs in our scenario, can be searched for at existing Large Hadron Collider (LHC) detectors. We here focus on a search for displaced showers in the CMS muon system Tumasyan et al. (2021), specifically the muon endcap detectors furthest from the interaction point. Such displaced showers originate when the LLP decays to photons, electrons, taus or quarks. The original search was for a generic scalar LLP $S$, pair-produced via the SM Higgs, $pp\to h\to SS$ and decaying to bottom quarks, down quarks or $\tau$ leptons. It provides the most stringent limits for long LLP proper decay lengths of $6-40$ m and LLP masses of $7-40$ GeV. Similarly, ATLAS searched for hadronic LLP decays Aaboud et al. (2019); Aad et al. (2020, 2023b), with weaker limits. ATLAS search  Aad et al. (2023b) has also been reinterpreted for pair production of HNL Wang (2024). The CMS muon endcap search has been considered in Cottin et al. (2023) for sterile neutrinos. We instead reinterpret it in the $B-L$ gauge model. We especially focus on Higgs production and we find that it is sensitive to the canonical seesaw floor, i.e., for active-sterile mixing strengths required to generate the light neutrino masses. While the $B-L$ gauge model clearly makes more assumptions, and large enough Higgs mixing is required, our analysis shows for the first time that existing data from the LHC probes the origin of neutrino masses in a well motivated scenario.
This paper is organized as follows. In Sec. II, we compare the different RHN production and decay channels. Our reinterpretation of the CMS muon endcap search is discussed in Sec. III. We present our results in Sec. IV, demonstrating that existing data is probing the canonical seesaw floor. We conclude in Sec. V.
@@ II Right-handed Neutrino Production and Decay
The $B-L$ gauge model extends the SM gauge group by an Abelian gauge symmetry $U(1)_{B-L}$. The SM fermions carry their canonical $Q_{B-L}$ charges and the three RHNs $N_{i}$ have a $Q_{B-L}=-1$ but are otherwise sterile. The $B-L$ symmetry is spontaneously broken by the vacuum expectation value (VEV) of a SM singlet Higgs $\Phi$ with $Q_{B-L}=2$. This results in a heavy $Z^{\prime}$ gauge boson and heavy Majorana RHNs. The relevant features of the $B-L$ gauge model are summarized in Appendix A, where we also review constraints. To emphasize the principle of our approach we assume that a single RHN $N$ is light enough to be produced and that it mixes dominantly with a single lepton flavour $\ell$ at a time through the active-sterile mixing $V_{\ell N}$. As the SM background for muons is overwhelming in the muon chamber, we concentrate on $\ell=e,\tau$ Cottin et al. (2023).
We first compare the four resonant RHN production channels at the LHC, namely via the $Z^{\prime}$, SM $Z$, SM-like Higgs $h$ and $(B-L)$-like Higgs $\Phi$. We omit production via the SM charged and neutral currents as they are suppressed by the small active-sterile mixing $|V_{\ell N}|$. The RHNs instead decay to SM particles via these currents while contributions from $Z^{\prime}$, $h$ and $\Phi$ are negligible. The dominant channels are $N\to\ell W\to\ell q\bar{q},\ell\ell\nu$, $N\to\ell^{\pm}\ell^{\mp}\nu$. As we consider light RHNs, $m_{N}<m_{W,Z}$, they undergo 3-body decays with the proper decay length given in Eq. (1). At the canonical seesaw floor, the RHNs thus have proper decay lengths $L_{N}^{0}\approx\mathcal{O}(10)$ m.
The produced RHNs are boosted, increasing the observed decay length to $L_{N}=\beta\gamma L_{N}^{0}$, with an average boost factor of $\langle\beta\gamma\rangle\approx 1-10$. Such long decay lengths are best probed in detectors far away from the interaction point (IP). This motivates the use of the muon system as it forms the outer layer of the CMS detector at a distance of $\approx 8-13$ m from the IP.
@@ II.1 Gauge Portal
With quarks and RHNs charged under $B-L$, the most direct channel to produce RHNs at the LHC is through the $Z^{\prime}$ gauge portal, $pp\to Z^{\prime}\to NN$. As a benchmark, we will later use $m_{Z^{\prime}}=500$ GeV and the production cross section at the 13 TeV LHC is $\sigma(pp\to Z^{\prime}\to NN)=\sigma(pp\to Z^{\prime})\times\text{Br}(Z^{\prime}\to NN)$, with $\sigma(pp\to Z^{\prime})\approx 0.4~{}\text{fb}\times(g_{B-L}/10^{-3})^{2}$. The branching ratio is Br$(Z^{\prime}\to NN)\approx 1/13$ for $m_{N}\ll m_{Z^{\prime}}$, leading to a total cross section of
@@
The most stringent limit from ATLAS resonance searches Aad et al. (2019b) is $g_{B-L}\lesssim 0.005$ for $m_{Z^{\prime}}=500$ GeV.
The SM $Z$ can mix with the $Z^{\prime}$, inducing the SM gauge portal $pp\to Z\to NN$, controlled by the effective coupling $\beta=g_{B-L}\sin\theta_{ZZ^{\prime}}$. The SM cross section at the 13 TeV LHC is $\sigma(pp\to Z)\approx 5.8\times 10^{7}$ fb Aaboud et al. (2017), and the decay branching ratio is $\text{Br}(Z\to NN)\approx 0.48\times\beta^{2}$ for $m_{N}\ll m_{Z}$. This results in a total cross section of
$\sigma(pp\to Z\to NN)\approx 28~{}\text{fb}\times\left(\frac{\beta}{10^{-3}}\right)^{2}.$ (3)
Limits from atomic parity violation are of the order $\beta\lesssim 10^{-3}$ Dev et al. (2021).
@@ II.2 Higgs Portal
The SM Higgs and the $B-L$ Higgs also mix, described by the Higgs mixing parameter $\sin\alpha$. RHNs can then be produced via the SM Higgs, $pp\to h\to NN$, where we only consider gluon-gluon fusion; while vector boson fusion and associated Higgs production are also possible, they contribute at most 10% to the overall cross section CERN and the resulting final states are unlikely to pass our trigger requirements. The $pp\to h\to NN$ production cross section is then $\sigma(pp\to h\to NN)=\cos^{2}\alpha\times\sigma(pp\to h)_{\text{SM}}\times\text{Br}(h\to NN)$ Accomando et al. (2017); Deppisch et al. (2018). Here, $\sigma(pp\to h)_{\text{SM}}=44\pm 4$ pb ($50\pm 7$ pb) is the SM Higgs production cross section at the 13 TeV (14 TeV) LHC via gluon-gluon fusion CERN ; LHC . The branching ratio can be approximated as Accomando et al. (2017)
@@
with the SM Higgs decay width $\Gamma_{h}^{\text{SM}}\approx 4.1$ MeV de Florian et al. (2016). It is suppressed by $\sin\alpha$ and the RHN Yukawa coupling with the $B-L$ Higgs $\Phi$, $y_{N}=m_{N}/\langle\Phi\rangle=2g_{B-L}m_{N}/m_{Z^{\prime}}$. The cross section is maximized for $m_{N}=m_{h}/\sqrt{10}\approx 40$ GeV,
@@
Both the Higgs mixing $\sin\alpha$ and the $(B-L)$ VEV $\langle\Phi\rangle$ are experimentally constrained through direct and indirect Higgs probes and $Z^{\prime}$ searches, respectively. In our analysis, we choose $\langle\Phi\rangle=3.75$ TeV. This is above the current lower limit from EW precision tests, $\langle\Phi\rangle\gtrsim 3.5$ TeV Cacciari et al. (2012); Alcaraz et al. (2006).
Similarly, RHNs can be produced via the $B-L$ Higgs $\Phi$. This time, the SM quarks and gluons will couple to $\Phi$ with a suppression as $\sigma(pp\to\Phi\to NN)\approx\sin^{2}\alpha\times\sigma(pp\to\Phi)_{\text{SM}}\times\text{Br}(\Phi\to NN)$, where $\sigma(pp\to\Phi)_{\text{SM}}$ is the production cross section treating $\Phi$ as the SM Higgs but with mass $m_{\Phi}$. The $B-L$ Higgs decays to heavy SM fermions and boson pairs via mixing to the SM Higgs, and RHN pairs via the Yukawa coupling $y_{N}$. For $m_{\Phi}<2m_{W}$, the branching ratio to RHN pairs is
@@
where $\Gamma_{\Phi}^{\text{SM}}$ is the total decay width treating $\Phi$ as the SM Higgs but with mass $m_{\Phi}$, and the partial width is $\Gamma(\Phi\to NN)=\cos^{2}\alpha/(16\pi)m_{N}^{2}m_{\Phi}/\langle\Phi\rangle^{2}$ for $m_{N}\ll m_{\Phi}$ Liu et al. (2022c). This results in the 13 TeV LHC cross section
@@
for $m_{\Phi}<2m_{W}\approx 160$ GeV. For $m_{\Phi}\gtrsim 2m_{W,Z}\approx 200$ GeV, the decays $\Phi\to WW,ZZ$ dominate, giving
@@
The cross section decreases rapidly with larger $m_{\Phi}$ and we use $m_{\Phi}=150$ GeV as benchmark in the following.

### v2

@@ Abstract
We study the potential to probe the origin of neutrino masses, by searching for long-lived right-handed neutrinos (RHNs) $N$ in the $B-L$ model and in the RHN-extended Standard Model (SM) Effective Field Theory (EFT). Despite the small active-sterile mixing $|V_{\ell N}|^{2}$, RHNs are produced abundantly via SM and exotic Higgs production, as long as the Higgs mixing or EFT operator coefficient is sufficiently large. We reinterpret a search for displaced showers in the CMS muon system and we find that it is sensitive to parameter space at and below the seesaw floor, $|V_{\ell N}|^{2}\approx 10^{-12}$ ($\ell=e$, $\tau$) for $m_{N}\approx 40$ GeV. With existing data constraining such well-motivated scenarios of neutrino mass generation, we determine the projected sensitivity at the HL-LHC, motivating dedicated searches for long-lived RHNs with decay lengths $\approx 10$ m.
@@ I Introduction
The Standard Model (SM) does not incorporate the observed neutrino masses confirmed in neutrino oscillations [1, 2, 3, 4], with the most popular explanation being the seesaw mechanism [5] where heavy right-handed neutrinos (RHNs) are introduced. If RHN masses $m_{N}$ are of electroweak (EW) scale, the active-sterile mixing strength $V_{\ell N}$ required to generate the neutrino mass $m_{\nu}$ is $|V_{\ell N}|^{2}\sim m_{\nu}/m_{N}\lesssim 10^{-11}\times(40~{}\text{GeV}/m_{N})$. While the absolute light neutrino mass scale has not been determined yet, the KATRIN experiment limits the effective $\beta$ decay mass to $m_{\beta}<0.45$ eV at 90 % confidence level (CL) [6], whereas solar neutrino oscillations point toward a smallest scale $m_{\nu}>\sqrt{\Delta m^{2}_{\text{sol}}}=9\times 10^{-3}$ eV [7]. This makes RHNs long-lived and given the displaced signature, RHNs have received increased attention in phenomenological studies [8, 9, 10, 11, 12, 13, 14, 15, 16, 17], and numerous collider searches have been carried out [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30].
While suggestive, the seesaw mechanism does not explain the origin of RHN Majorana masses. This is addressed in models where lepton number is broken spontaneously, generating the RHN masses. This includes, for example, the $U(1)_{B-L}$ gauge model and extensions [31, 32, 33, 34], models based on Left-Right [35, 36, 37] and Pati-Salam [38] symmetry as well as certain Two-Higgs-Doublet models [39]. Such scenarios generically predict exotic gauge and scalar bosons that mix with the corresponding SM states, enabling new portals of RHN production. We here consider the $B-L$ (baryon - lepton number) model [31, 32] as a prominent and minimal example, where masses are generated through spontaneous breaking of a $U(1)_{B-L}$ gauge symmetry. The model predicts three RHNs $N_{i}$, an exotic gauge boson $Z^{\prime}$ and a $B-L$-breaking Higgs $\Phi$. The RHNs may also explain the baryon asymmetry via leptogenesis [40, 41, 42, 17] and the $B-L$ model can be conformally invariant, not only explaining why $B-L$ is broken near the TeV scale and inducing EW symmetry breaking [43, 44], but also predicting primordial black holes as dark matter [45]. The low energy effects of this and similar models, at SM scales and below, can also be described within the RHN-extended extended SM Effective Field Theory ($\nu_{R}$) framework [46, 47, 48, 49, 50].
RHNs in the $B-L$ model not only interact with the SM particles via the active-sterile neutrino mixing $V_{\ell N}$, but also through the $B-L$ gauge and Higgs portals. Thus, RHNs can be produced via the SM Higgs, $B-L$ Higgs, SM $Z$ and $B-L$ $Z^{\prime}$. It is well known that enabling these portals allows testing small active-sterile mixing strengths [51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 47, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73], as the RHNs still decay as $N\to\nu_{\ell}h,\nu_{\ell}Z^{*},\ell W^{*}$, controlled by $|V_{\ell N}|$ ($\ell=e,\mu,\tau$). This leads to macroscopic proper decay lengths ($m_{N}<m_{W,Z}$[74],
@@
Long-lived particles (LLPs), such as RHNs, can be searched for at Large Hadron Collider (LHC) detectors. We here focus on a search for displaced showers in the CMS muon endcap [75] for a scalar LLP $S$, pair-produced via the Higgs, $pp\to h\to SS$, and decaying to quarks or $\tau$ leptons. It provides the most stringent limit for LLP proper decay lengths $6-40$ m and LLP masses $7-40$ GeV. ATLAS searched for hadronic LLP decays [76, 77, 78] but with weaker limits. The CMS muon endcap search has been considered in [79] for sterile neutrinos but we reinterpret it in the $B-L$ model and $\nu_{R}$, focusing on Higgs production where we find that it is sensitive to the seesaw floor, i.e., for active-sterile mixing required to generate light neutrino masses. Our analysis shows for the first time that existing data from the LHC probes the origin of neutrino masses in well motivated scenarios.
@@ II Right-handed Neutrino Production and Decay
The $B-L$ model extends the SM by an Abelian gauge symmetry $U(1)_{B-L}$. It is spontaneously broken by the vacuum expectation value (VEV) of a SM-singlet Higgs $\Phi$, resulting in a heavy $Z^{\prime}$ gauge boson and heavy Majorana RHNs. The relevant features of and constraints on the $B-L$ model are summarized in Appendix A. For simplicity, we assume that a single RHN $N$ is light enough to be produced and that it mixes with a single lepton flavour $\ell$ at a time through $V_{\ell N}$. We concentrate on $\ell=e,\tau$ due to the search discussed in the following section: First, RHN LLPs decaying to muons rarely produce a particle shower required by the search and second, the search vetoes events with muons too close to a jet [75]. We therefore expect no sensitivity to muon flavour.
We first compare the four resonant RHN production channels at the LHC, namely via the $Z^{\prime}$, SM $Z$, SM-like Higgs $h$ and $(B-L)$-like Higgs $\Phi$. We omit production via the SM charged and neutral currents as they are suppressed by the small active-sterile mixing $|V_{\ell N}|$. The RHNs instead decay to SM particles via these currents with decay lengths $L_{N}^{0}\approx 10$ m at the seesaw floor, see Eq. (1). Produced from electroweak-scale or heavier states, the RHNs are boosted, increasing the observed decay length to $L_{N}=\beta\gamma L_{N}^{0}$, with an average boost factor $\langle\beta\gamma\rangle\approx 1-10$. Such long decay lengths are best probed in detectors far away from the interaction point (IP), namely the muon system, forming the outer layer of the CMS detector at a distance of $\approx 8-13$ m from the IP.
@@ II.1 Gauge Portal
With quarks and RHNs charged under $B-L$, RHNs are produced at the LHC through the $Z^{\prime}$ gauge portal, $pp\to Z^{\prime}\to NN$. As a benchmark, we use $m_{Z^{\prime}}=500$ GeV and the cross section at the 13 TeV LHC is $\sigma(pp\to Z^{\prime}\to NN)\approx\sigma(pp\to Z^{\prime})\times\text{Br}(Z^{\prime}\to NN)$, with $\sigma(pp\to Z^{\prime})\approx 0.4~{}\text{fb}\times(g_{B-L}/10^{-3})^{2}$. The branching ratio is Br$(Z^{\prime}\to NN)\approx 1/13$ for $m_{N}\ll m_{Z^{\prime}}$, leading to
@@
The most stringent limit from ATLAS resonance searches [80] is $g_{B-L}\lesssim 5\times 10^{-3}$ for $m_{Z^{\prime}}=500$ GeV.
The SM $Z$ can mix with the $Z^{\prime}$, inducing the SM gauge portal $pp\to Z\to NN$, controlled by the effective coupling $\beta=g_{B-L}\sin\theta_{ZZ^{\prime}}$. The SM cross section at the 13 TeV LHC is $\sigma(pp\to Z)\approx 5.8\times 10^{7}$ fb [81], and the branching ratio is $\text{Br}(Z\to NN)\approx 0.48\times\beta^{2}$ for $m_{N}\ll m_{Z}$. This results in
$\sigma(pp\to Z\to NN)\approx 28~{}\text{fb}\times\left(\frac{\beta}{10^{-3}}\right)^{2},$ (3)
consistent with atomic parity violation limits [82].
@@ II.2 Higgs Portal
The SM Higgs and the $B-L$ Higgs also mix, through the Higgs mixing parameter $\sin\alpha$. RHNs can then be produced via the SM Higgs, $pp\to h\to NN$, where we only consider gluon-gluon fusion; while vector boson fusion and associated Higgs production are also possible, they contribute at most 10% to the overall cross section [83] and the resulting final states are too soft to pass the missing transverse energy requirement $E_{T}^{\text{miss}}>200$ GeV, discussed in Sec. III. The $pp\to h\to NN$ production cross section is $\sigma(pp\to h\to NN)\approx\cos^{2}\alpha\times\sigma(pp\to h)_{\text{SM}}\times\text{Br}(h\to NN)$ [66, 64]. Here, $\sigma(pp\to h)_{\text{SM}}=44\pm 4$ pb ($50\pm 7$ pb) is the SM Higgs production cross section at the 13 TeV (14 TeV) LHC via gluon-gluon fusion [83, 84]. The branching ratio can be approximated as [66]
@@
with the SM Higgs width $\Gamma_{h}^{\text{SM}}\approx 4.1$ MeV [85]. It is suppressed by $\sin\alpha$ and the RHN Yukawa coupling with the $B-L$ Higgs $\Phi$, $y_{N}=m_{N}/\langle\Phi\rangle=2g_{B-L}m_{N}/m_{Z^{\prime}}$. The cross section is maximized for $m_{N}=m_{h}/\sqrt{10}\approx 40$ GeV,
@@
Both the Higgs mixing $\sin\alpha$ and the $B-L$ VEV $\langle\Phi\rangle$ are experimentally constrained through direct and indirect Higgs probes and $Z^{\prime}$ searches. In our analysis, we choose $\langle\Phi\rangle=3.75$ TeV, satisfying the current limit from EW precision tests, $\langle\Phi\rangle\gtrsim 3.5$ TeV [86, 87].
Similarly, RHNs can be produced via $\Phi$. The SM quarks and gluons couple to $\Phi$ with a suppression as $\sigma(pp\to\Phi\to NN)\approx\sin^{2}\alpha\times\sigma(pp\to\Phi)_{\text{SM}}\times\text{Br}(\Phi\to NN)$, where $\sigma(pp\to\Phi)_{\text{SM}}$ is the cross section treating $\Phi$ as the SM Higgs but with mass $m_{\Phi}$. $\Phi$ decays to heavy SM fermions and boson pairs via mixing to the SM Higgs, and RHN pairs via the Yukawa coupling $y_{N}$. For $m_{\Phi}<2m_{W}$, the branching ratio is
@@
where $\Gamma_{\Phi}^{\text{SM}}$ is the decay width treating $\Phi$ as the SM Higgs but with mass $m_{\Phi}$, and the partial width is $\Gamma(\Phi\to NN)=\cos^{2}\alpha/(16\pi)m_{N}^{2}m_{\Phi}/\langle\Phi\rangle^{2}$ for $m_{N}\ll m_{\Phi}$ [70]. This results in the 13 TeV LHC cross section
@@
for $m_{\Phi}<2m_{W}\approx 150$ GeV. For $m_{\Phi}\gtrsim 2m_{W,Z}\approx 200$ GeV, the decays $\Phi\to WW,ZZ$ dominate,
@@
The cross section decreases rapidly with larger $m_{\Phi}$ and we use $m_{\Phi}=150$ GeV as benchmark.